what is the answer for 26=3(2x+1)+5




Answer by addingup(3677)

I call it the "do no harm" rule of algebra, whatever you do to one side you also do to the other side. In this case:
:
26 = 3(2x+1)+5
26-5 = 3(2x+1)+5-5
21 = 3(2x+1) Resolve on right:
21 = 3*2x+3*1
21 = 6x+3
21-3 = 6x+3-3
18 = 6x
18/6 = 6x/6
3 = x
or
x = 3